OpenStudy (anonymous):

The sum of the digits of a two digit number is 12. The number is 7 times the sum of it's digits. What's the number?

OpenStudy (oaktree):

If the number is 10a + b, then:

OpenStudy (oaktree):

a+b=12, and 10a+b=7(a+b)

OpenStudy (oaktree):

So, 10a+b

OpenStudy (anonymous):

i got it from there. Thank you!

OpenStudy (oaktree):

Is equal to 84.